OpenStudy (anonymous):

the line spectrum of lithium has a red line at 670.8 nm. calculate the energy of a photon with this wavelength

OpenStudy (anonymous):

E = ħc / λ ħ = plancks constant = 6.626x10^-34 Js c = speed of light = 2.999x10^8 m/s λ = wavelength of light = 670.8x10^-9 m E = (6.626x10^-34 Js) x (2.999x10^8 m/s) x (1 / 670.8x10^-9 m) E = 2.962x10^-19 J

OpenStudy (anonymous):

3x10^8 / (670.8 * 10^-9) =4.47x10^14 Hz 4.47x10^14 Hz multiplied by plank's constant = 2.9634x10^-19

OpenStudy (anonymous):

either of these answers will work for you

OpenStudy (preetha):

Good job Helter.